If the resistor are in parallel,prove also that the equivalent resistance is R where

1/R=1/R1 +1/R2 +1/R3+...1/RN


1
Expert's answer
2021-03-16T11:38:23-0400

Explanations & Calculations


  • For the array of resistors, the total current by it can be written as follows as the total current is the sum of currents through each branch: resistor.

"\qquad\qquad
\begin{aligned}
\small I&=\small i_1+i_2+i_3+i_4+.....+i_n
\end{aligned}"

  • As the potential drop across each resistor is equal in a parallel arrangement, individual currents from the above equation can be replaced by "i=\frac{V}{R}",

"\qquad\qquad
\begin{aligned}
\small I&=\small \frac{V}{R_1}+\frac{V}{R_2}+\frac{V}{R_3}+\frac{V}{R_4}+.....+\frac{V}{R_n}
\end{aligned}"

  • An equivalent resistance "\small R" to those array of resistors is such, that it conducts the same total current "\small I" having the same potential drop "\small V" across it.
  • Then,

"\qquad\qquad
\begin{aligned}
\small I&=\small\frac{V}{R}
\end{aligned}"

  • Then replacing the total current "\small I" from both the equations we get

"\qquad\qquad
\begin{aligned}
\small \frac{V}{R}&=\small \frac{V}{R_1}+\frac{V}{R_2}+\frac{V}{R_3}+\frac{V}{R_4}+....+\frac{V}{R_n}\\
\small \frac{1}{R}&=\small \frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}+\frac{1}{R_4}+.......+\frac{1}{R_n}\cdots(\,\cancel{\small V}\,)
\end{aligned}"